OpenStudy (anonymous):

Are these right? 0_0 Write the following number 3 different ways. The number is 19 19/1 19.000 19%

OpenStudy (dan815):

not 19% lol

OpenStudy (dan815):

19% = 19/100 = 0.19

OpenStudy (anonymous):

OK, someone told me it was .19. Can you explain that? :)

OpenStudy (dan815):

you can write 19 in all kinds of ways as long as it just simplifies to 19 like 38/2

hero (hero):

I'd go with 38/2 sqrt(361) and 1900%

OpenStudy (dan815):

or 20-1

OpenStudy (anonymous):

Thanks! :) I think I got it. 19 as a percent would be 0.19? Or .19?

OpenStudy (marissalovescats):

0.19 and .19 is the same thing, you just don't have the zero there. Just like 0.19 and 0.190 is the same thing.

hero (hero):

\[19 = 19 \times 1 = 19 \times \frac{100}{100} = 1900\left( \frac{1}{100}\right)= 1900\%\]